RHIC Startup:
Several 56x56 stores were lost last night when there was a Beam Permit pulled followed several seconds later by a Quench Link Interlock, not a quench. Attempts to store a beam at intensities of 1.06E11/1.09E9, 0.98E11/, 0.86E11/0.6E9 .
Wolfram and Peggy will begin by working on aligning the abort gaps in Blue and Yellow.
MikeB found that at the frequency jump, two of the storage cavities suffer a drop out in voltage. May request a high intensity ramp to see if it gets worse, Also would like to try the stochastic cooling in Yellow before tomorrow's maintenance day. Want to insert the three tanks and scan the beam to see if the beam is centered in the BPMs. And stochastic cooling is ready to go. Can be accomplished with a 6x6 ramp. Need about an hour, immediately following Wolfram and Peggy. Assuming it works well, it could be turned on tonight. Check aperture with 6x6 ramp.
If BBQ is working, both Blue and Yellow tunes should be flat. Currently only the Blue is flat. Why can't we make Yellow flat? Coupling data is good enough to do a feedback ramp. Need about 3 hours to do this.
Mei wants to measure the beta* at store in Yellow, since yesterday's measurement was compromised by the rebucketing. Will take about 20 minutes at injection. Mei also wants to do orbit measurements in Blue at injection.
Al wants to get the Yellow feedback ramp and replay going.
Do orbit, coupling and chromaticity corrections.
Wolfram wants to measure the tune shift from the front to the back of the bunch trains. .
Wolfram thinks it may be important to do some scrubbing. Vacuum wants to be notified the day before, and requires 3 hours set-up. Possibly do scrubbing on Thursday.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email:sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |